# TableDART: Dynamic Adaptive Multi-Modal Routing for Table Understanding

### v1

@@ Abstract
Modeling semantic and structural information from tabular data remains a core challenge for effective table understanding. Existing Table-as-Text approaches flatten tables for large language models (LLMs), but lose crucial structural cues, while Table-as-Image methods preserve structure yet struggle with fine-grained semantics. Recent Table-as-Multimodality strategies attempt to combine textual and visual views, but they (1) statically process both modalities for every query-table pair within a large multimodal LLMs (MLLMs), inevitably introducing redundancy and even conflicts, and (2) depend on costly fine-tuning of MLLMs. In light of this, we propose TableDART, a training-efficient framework that integrates multimodal views by reusing pretrained single-modality models. TableDART introduces a lightweight 2.59M-parameter MLP gating network that dynamically selects the optimal path (either Text-only, Image-only, or Fusion) for each table–query pair, effectively reducing redundancy and conflicts from both modalities. In addition, we propose a novel agent to mediate cross-modal knowledge integration by analyzing outputs from text- and image-based models, either selecting the best result or synthesizing a new answer through reasoning. This design avoids the prohibitive costs of full MLLM fine-tuning. Extensive experiments on seven benchmarks show that TableDART establishes new state-of-the-art performance among open-source models, surpassing the strongest baseline by an average of 4.02%. The code is available at: https://anonymous.4open.science/r/TableDART-C52B.
@@ 1 Introduction
Tabular data is one of the most ubiquitous formats in the real world, with applications spanning finance (Chen et al., 2021; Zhu et al., 2021; 2024b), healthcare (Shi et al., 2024; Wang et al., 2024), and many other domains that rely on relational databases and spreadsheets (Fey et al., 2024). Despite its prevalence, effectively modeling tabular data remains a long-standing challenge, due to its unique characteristics such as heterogeneity, permutation invariance, and hierarchical structure (Fang et al., 2024; Borisov et al., 2024; Chen et al., 2023).
Existing approaches to tabular data understanding can be broadly categorized into two paradigms. The first is the “Table-as-Text” paradigm, where tables are linearized into sequences and processed by large language models (LLMs) (Zha et al., 2023; Su et al., 2024). While effective, this approach is sensitive to serialization choices (Sui et al., 2024) and inevitably loses critical structural information (Deng et al., 2024). The second paradigm treats “Table-as-Image”, where table screenshots are processed with vision models. This approach better preserves structural information but struggles to capture fine-grained and aligned semantic meaning (Deng et al., 2024).
To combine the complementary strengths of these paradigms, recent research has moved toward “Table-as-Multimodality”, where tables are represented in both textual and visual forms and processed via fine-tuned multimodal LLMs (MLLMs) (Liu et al., 2025). While promising, this MLLM-based paradigm suffers from two key limitations that hinder both performance and practicality. First, regardless of the context, they mandate both modality views for every table-query pair. However, not all queries benefit from multiple views, as integrating multiple views often introduces redundancy and, in some cases, conflicting signals. For example, textual linearization may inadvertently impose row-order sensitivity (Fang et al., 2024), while tables in their original or visual form remain permutation-invariant (Wu et al., 2025). Such redundancy and conflicts mislead the MLLM’s table understanding, resulting in suboptimal performance. Second, existing approaches typically require fine-tuning large MLLMs, which is computationally prohibitive to scale, even with efficient fine-tuning strategies (Hu et al., 2022).
@@ 2 Related Work
Unimodal Approaches for Table Understanding. A dominant paradigm for applying LLMs to tabular data is the Table-as-Text approach, which serializes tables into linear text sequences. Early work such as TaPas (Herzig et al., 2020) demonstrated its effectiveness, inspiring specialized tabular models like TableLlama (Zhang et al., 2024) and TableGPT2 (Su et al., 2024), which achieve state-of-the-art results for table understanding. However, these models remain constrained by their input modality, as they cannot capture the rich visual and structural semantics inherent in the original table format (Deng et al., 2024). To overcome this limitation, the Table-as-Image paradigm has emerged, leveraging Multi-modal Large Language Models (MLLMs) to directly process table images. Pioneering works such as Table-LLaVA (Zheng et al., 2024) and generalist MLLMs like Ovis2 (Lu et al., 2024) showcase the potential of modeling visual layouts and structural cues. Yet, this paradigm is not a universal solution, as its effectiveness diminishes for tasks where reasoning relies heavily on text-centric pretraining knowledge (Deng et al., 2024).
@@ 3.2 Encoding Table-as-Multimodality
Given a table $\mathcal{T}$ and a query $q$, TableDART leverages existing models to encode them from complementary modality perspectives, thereby capturing both semantic and structural information. This is achieved through a parallel feature extraction pipeline with three concurrent streams: (1) Table-as-Text encoding: The table $\mathcal{T}$ is serialized into text and processed by the encoder $\mathcal{E}_{t}$ of the Table-as-Text model $\mathcal{M}_{t}$, yielding a feature vector $\mathbf{e}_{t}$. (2) Table-as-Image encoding: A screenshot of $\mathcal{T}$ is fed into the encoder $\mathcal{E}_{v}$ of the Table-as-Image model $\mathcal{M}_{v}$, producing $\mathbf{e}_{v}$. (3) Query encoding: The query $q$ is transformed into an embedding $\mathbf{e}_{q}$ using a text embedding model $\mathcal{E}_{q}$. The resulting feature vectors are concatenated to form a multimodal representation of the table-query pair:
$\mathbf{e}_{t}=\mathcal{E}_{t}(Serialize(\mathcal{T}))$ (1)
$\mathbf{e}_{v}=\mathcal{E}_{v}(Screenshot(\mathcal{T}))$
@@
Here, $\mathbf{x}$ serves as the multimodal joint representation by concatenating the three embeddings, which is subsequently passed into the gating network to determine the path of dynamic inference. Note that $\mathcal{E}_{t}$ and $\mathcal{E}_{v}$ only activate a small part of the corresponding expert models, constituting a minor computational overhead (i.e., 7.15% of the Table-as-Text model’s and 7.63% of the Table-as-Image model’s total parameters as shown in Appendix A.3.1).
@@ 3.3 Gating Network and Policy Training
$\mathbf{z}=\mathcal{G}(\mathbf{x})$ (2)
@@
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{task}}+\lambda\mathcal{L}_{\text{resource}}$ (3)
@@
The task loss $\mathcal{L}_{\text{task}}$ encourages the gating network to prioritize paths with strong empirical performance. We compute a target distribution for each training instance by first evaluating the success or failure of each path to obtain a vector of binary scores, and then normalizing this vector into a soft probability distribution using a temperature-controlled softmax. The gating network’s predicted distribution is then guided to align with this target by minimizing the KL divergence:
$\mathcal{L}_{\text{task}}=\text{KL}(\text{softmax}(\mathbf{s}/\tau)\parallel\text{softmax}(\mathbf{z}/\tau_{g}))$ (4)
where $\mathbf{z}$ represents the raw logits from the gating network, $\mathbf{s}$ is the vector of accuracy scores for each inference path, and $\tau_{g},\tau$ are temperature parameters.
@@
$\mathcal{L}_{\text{resource}}=\text{softmax}(\mathbf{z}/\tau_{g})^{T}\mathbf{c}$ (5)
@@ 3.4 Adaptive Inference
TableDART supports three inference paths: Text-only, Image-only, and Fusion. The first two directly reuse existing single-modality expert models: the Table-as-Text model $\mathcal{M}_{t}$ and the Table-as-Image model $\mathcal{M}_{v}$. If the gating network selects either, TableDART simply continues the inference process from the intermediate representation of the chosen model. For example, when the gate selects the Text-only path, TableDART forwards $\mathcal{M}_{t}$’s remaining layers using the intermediate encoding $\mathbf{e}_{t}$ to produce the final output.
For the Fusion path, we design a novel, training-free LLM-based Fusion agent that synthesizes the final answer by reasoning over the outputs of the two single-modality experts, $\mathcal{M}_{t}$ and $\mathcal{M}_{v}$. Specifically, when the gating network selects the Fusion path, both $\mathcal{M}_{t}$ and $\mathcal{M}_{v}$ are at first executed to generate results $r_{t}$ and $r_{v}$, along with their auxiliary outputs $a_{t}$ and $a_{v}$. These, together with the original table $\mathcal{T}$, are passed to the Fusion agent. The Fusion agent operates in two possible roles depending on its reasoning process: (1) Arbitrator: If $\mathcal{M}_{t}$ and $\mathcal{M}_{v}$ produce conflicting results, the agent resolves the disagreement by selecting the more reliable answer according to its confidence. (2) Rescuer: If the agent believes both models provide uncertain or low-confidence outputs, it synthesizes a new, more accurate answer by jointly reasoning over their partial evidence. A detailed description of the Fusion agent’s implementation and prompt design is provided in Appendix A.4.
@@ 4.1 Experimental Setup
Evaluation Datasets and Metrics. We evaluate TableDART on seven diverse benchmarks across two primary tasks: Table Question Answering (TQA)and Table Fact Verification (TFV). For TQA, we use five benchmarks: WTQ (Pasupat & Liang, 2015), TABMWP (Lu et al., 2023), TAT-QA (Zhu et al., 2021), HiTab (Cheng et al., 2022), and FeTaQA (Nan et al., 2022). For TFV, we use TabFact (Chen et al., 2020) and InfoTabs (Gupta et al., 2020). Following the established protocol (Zheng et al., 2024; Liu et al., 2025), we use Accuracy for WTQ, TABMWP, TAT-QA, HiTab, TabFact, InfoTabs, while using BLEU score for FeTaQA as it is a generative free-form response task. Detailed statistics and further experimental setup details are provided in Appendix A.5.
Baselines. We compare TableDART against a comprehensive set of baselines, including (1) Table-as-Text models: Llama-2-7B (Touvron et al., 2023), Llama3-Instruct-8B (Dubey et al., 2024), TableLlama-7B (Zhang et al., 2024), and TableGPT2-7B (Su et al., 2024); (2) Table-as-Image models: MiniGPT-4-7B (Zhu et al., 2024a), mPLUG-Owl-7B (Ye et al., 2023), mPLUG-Owl2-7B (Ye et al., 2024), LLaVA v1.5-7B (Liu et al., 2024), Table-LLaVA-7B (Zheng et al., 2024), Qwen-VL-7B Bai et al. (2023), InternLM-XComposer2-7B (Zhang et al., 2023), Monkey-7B (Li et al., 2024), MiniCPM-V-2.6-8B (Yao et al., 2024), and Ovis2-8B (Lu et al., 2024); (3) Table-as-Multimodality models: HIPPO (Liu et al., 2025) and Google Gemini 2.0 Flash (Comanici et al., 2025).
We employ TableGPT2-7B and Ovis2-8B as TableDART’s single-modality experts and use Google Gemini 2.0 Flash to implement the Fusion agent. Comparing TableDART against these base models directly demonstrates the effectiveness of our framework. A detailed rationale for model selection is provided in Appendix A.5. Notably, TableDART is a general framework that can be seamlessly integrated with a wide range of existing LLMs, VLMs and MLLMs .
@@ 4.2 Effectiveness Analysis
TableDART 70.58 84.54 62.05 74.37 36.11 81.37 76.22 74.86
State-of-the-Art Performance. Table 4.2 shows the comparison results of TableDART with baselines. From the results, we can observe that: (1) Compared to all single-modality based methods, TableDART consistently achieves superior or highly competitive results, including surpassing its constituent models like TableGPT2-7B and Ovis2-8B, demonstrating that the framework successfully combines their strengths to achieve a result neither could reach alone. (2) Compared to multi-modality based methods, TableDART’s dynamic routing mechanism proves more effective than the MLLM-based HIPPO, outperforming it on five of the seven benchmarks. (3) TableDART outperforms the powerful Fusion agent model (Google Gemini 2.0 Flash), confirming that the performance gains are driven by the intelligent routing mechanism, not just the capacity of the Fusion agent’s backbone. Furthermore, the ‘Average’ column summarizes the average performance, showing that TableDART achieves the strongest results among all baselines, surpassing the best multi-modality tabular model, HIPPO-8B, by a decisive +4.02% in average accuracy and +2.93 points in BLEU score.

### v2

@@ Abstract
Modeling semantic and structural information from tabular data remains a core challenge for effective table understanding. Existing Table-as-Text approaches flatten tables for large language models (LLMs), but lose crucial structural cues, while Table-as-Image methods preserve structure yet struggle with precise semantics. Recent Table-as-Multimodality strategies attempt to combine textual and visual views, but they (1) statically process both modalities for every query-table pair within large multimodal LLMs (MLLMs), inevitably introducing redundancy and even conflicts, and (2) depend on costly fine-tuning of MLLMs. In light of this, we propose TableDART, a training-efficient framework that integrates multimodal views by reusing pretrained single-modality models. TableDART introduces a lightweight 2.59M-parameter MLP gating network that dynamically selects the optimal path (Text-only, Image-only, or Fusion) for each table–query pair, reducing redundancy and avoiding conflicts that arise when textual and visual views of the same table provide inconsistent cues. By routing to the most appropriate view, our framework improves both accuracy and efficiency. In addition, we propose a novel agent to mediate cross-modal knowledge integration by analyzing outputs from text- and image-based models, either selecting the best result or synthesizing a new answer through reasoning. This design avoids the prohibitive costs of full MLLM fine-tuning. Extensive experiments on seven benchmarks show that TableDART establishes new state-of-the-art performance among open-source models, surpassing the strongest baseline by an average of 4.02%. The code is available at: https://github.com/xiaobo-xing/TableDART.
[…]
@@ 1 Introduction
Tabular data is one of the most ubiquitous formats in the real world, with applications spanning finance (Chen et al., 2021; Zhu et al., 2021; 2024b), healthcare (Shi et al., 2024; Wang et al., 2024), and many other domains that rely on relational databases and spreadsheets (Fey et al., 2024; Gao et al., 2026). Despite its prevalence, effectively modeling tabular data remains a long-standing challenge, due to its unique characteristics such as heterogeneity of diverse data types across columns, permutation invariance regarding row and column order, and hierarchical structure of multi-level or nested headers and indices (Fang et al., 2024; Borisov et al., 2024; Chen et al., 2023).
Existing approaches to tabular data understanding can be broadly categorized into two paradigms. The first is the “Table-as-Text” paradigm, where tables are linearized into sequences and processed by large language models (LLMs) (Zha et al., 2023; Su et al., 2024). While effective, this approach is sensitive to serialization choices (Sui et al., 2024) and inevitably loses critical structural information (Deng et al., 2024). The second paradigm treats “Table-as-Image”, where table screenshots are processed with vision models. This approach better preserves structural information but struggles to capture precise and aligned semantic meaning (Deng et al., 2024).
To combine the complementary strengths of these paradigms, recent research has moved toward “Table-as-Multimodality”, where tables are represented in both textual and visual forms and processed via fine-tuned multimodal LLMs (MLLMs) (Liu et al., 2025). While promising, this MLLM-based paradigm suffers from two key limitations that hinder both performance and practicality. First, regardless of the context, they mandate both modality views for every table-query pair. However, not all queries benefit from multiple views, as integrating multiple views often introduces redundancy and, in some cases, conflicting signals. For example, textual linearization may inadvertently impose row-order sensitivity (Fang et al., 2024), while tables in their original or visual form remain permutation-invariant (Wu et al., 2025). Such redundancy and conflicts mislead the MLLM’s table understanding, resulting in suboptimal performance. Second, existing approaches typically rely on heavy MLLMs, which remain computationally prohibitive to adapt at scale even with parameter-efficient fine-tuning strategies (Hu et al., 2022), limiting their practicality under constrained training budgets and deployment pipelines (Yin et al., 2025).
@@ 2 Related Work
Unimodal Approaches for Table Understanding. A dominant paradigm for applying LLMs to tabular data is the Table-as-Text approach, which serializes tables into linear text sequences. Early work such as TaPas (Herzig et al., 2020) demonstrated its effectiveness, inspiring specialized tabular models like TableLlama (Zhang et al., 2024) and TableGPT2 (Su et al., 2024), which achieve state-of-the-art results for table understanding. However, these models remain constrained by their input modality, as they cannot capture the rich visual and structural semantics inherent in the original table format (Deng et al., 2024). To overcome this limitation, the Table-as-Image paradigm has emerged, leveraging Multi-modal Large Language Models (MLLMs) to directly process table images. Pioneering works such as Table-LLaVA (Zheng et al., 2024), TabPedia (Zhao et al., 2024), SynTab-LLaVA (Zhou et al., 2025), and generalist MLLMs like Ovis2 (Lu et al., 2024) and Qwen2.5-VL (Bai et al., 2025) showcase the potential of modeling visual layouts and structural cues, aligning with broader trends in vision-centric text understanding (Yuan et al., 2026; Qiao et al., 2026). Yet, this paradigm is not a universal solution, as its effectiveness diminishes for tasks where reasoning relies heavily on text-centric pretraining knowledge (Deng et al., 2024).
@@ 3.2 Encoding Table-as-Multimodality
Given a raw table $\mathcal{T}$ and a query $q$, TableDART leverages existing models to encode them from complementary modality perspectives, thereby capturing both semantic and structural information. This is achieved through a parallel feature extraction pipeline with three concurrent streams: (1) Table-as-Text encoding: The raw table $\mathcal{T}$ is serialized into text and processed by the encoder $\mathcal{E}_{t}$ of the Table-as-Text model $\mathcal{M}_{t}$, yielding a feature vector $\mathbf{e}_{t}$. (2) Table-as-Image encoding: A screenshot of $\mathcal{T}$ is fed into the encoder $\mathcal{E}_{v}$ of the Table-as-Image model $\mathcal{M}_{v}$, producing $\mathbf{e}_{v}$. (3) Query encoding: The query $q$ is transformed into an embedding $\mathbf{e}_{q}$ using a text embedding model $\mathcal{E}_{q}$. We apply modality-specific pooling to convert encoder outputs into fixed-dimensional vectors, which are concatenated to form a unified multimodal representation for the gating network:
$\mathbf{e}_{t}=\mathcal{E}_{t}(\text{Serialize}(\mathcal{T}))$ (1)
$\mathbf{e}_{v}=\mathcal{E}_{v}(\text{Screenshot}(\mathcal{T}))$
@@
Here, $\mathbf{x}$ serves as the multimodal joint representation by concatenating the three embeddings, which is subsequently passed into the gating network to determine the path of dynamic inference. Note that $\mathcal{E}_{t}$ and $\mathcal{E}_{v}$ only activate a small part of the corresponding expert models, constituting a minor computational overhead (i.e., 7.15% of the Table-as-Text model’s and 7.63% of the Table-as-Image model’s total parameters). Detailed analyses of the embedding construction and parameter breakdown are provided in Appendix A.3.1.
@@ 3.3 Gating Network and Policy Training
$\mathbf{z}=\mathcal{G}(\mathbf{x}).$ (2)
@@
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{task}}+\lambda\mathcal{L}_{\text{resource}},$ (3)
@@
The task loss $\mathcal{L}_{\text{task}}$ encourages the gating network to prioritize paths with strong empirical performance. For each training instance, we pre-compute a vector of binary scores $\mathbf{s}=(s_{1},s_{2},s_{3})\in\{0,1\}^{3}$, where each component $s_{k}$ is 1 if the $k$-th inference path produces a correct answer, and 0 otherwise. This method naturally handles instances where multiple paths are correct, as each successful path receives a score of 1. To form the training signal, this score vector is converted into a soft target probability distribution using a temperature-controlled softmax. Similarly, the gating network’s raw logits $\mathbf{z}$ are converted into a predicted distribution. The task loss then minimizes the KL divergence between these two distributions:
$\mathcal{L}_{\text{task}}=\text{KL}(\text{softmax}(\mathbf{s}/\tau)\parallel\text{softmax}(\mathbf{z}/\tau_{g})),$ (4)
where $\tau_{g},\tau$ are temperature parameters. We use the standard forward KL divergence, which encourages the gating network to assign probability to all empirically successful paths.
@@
$\mathcal{L}_{\text{resource}}=\text{softmax}(\mathbf{z}/\tau_{g})^{T}\mathbf{c},$ (5)
@@ 3.4 Adaptive Inference
The adaptive inference process begins with a deterministic routing decision guided by the trained gating network. For each query-table pair, the network produces a vector of raw logit scores, and the framework selects the single, optimal path corresponding to the highest score for execution. Once the path is selected, inference proceeds accordingly. TableDART supports three inference paths: Text-only, Image-only, and Fusion. The first two directly reuse existing single-modality expert models: the Table-as-Text model $\mathcal{M}_{t}$ and the Table-as-Image model $\mathcal{M}_{v}$. If the gating network selects either, TableDART simply continues the inference process from the intermediate representation of the chosen model. For example, when the gate selects the Text-only path, TableDART forwards $\mathcal{M}_{t}$’s remaining layers using the intermediate encoding $\mathbf{e}_{t}$ to produce the final output.
For the Fusion path, we design a novel, training-free LLM-based Fusion agent that synthesizes the final answer by reasoning over the outputs of the two single-modality experts, $\mathcal{M}_{t}$ and $\mathcal{M}_{v}$. Specifically, when the gating network selects the Fusion path, both $\mathcal{M}_{t}$ and $\mathcal{M}_{v}$ are at first executed to generate results $r_{t}$ and $r_{v}$, along with their auxiliary outputs $a_{t}$ and $a_{v}$. These, together with the original table $\mathcal{T}$, are passed to the Fusion agent. The Fusion agent operates in two possible roles depending on its reasoning process: (1) Arbitrator: If $\mathcal{M}_{t}$ and $\mathcal{M}_{v}$ produce conflicting results, the agent resolves the disagreement, similar in spirit to aggregating a consensus from multiple sources (Hung et al., 2018), by selecting the more reliable answer according to its confidence. (2) Rescuer: If the agent believes both models provide uncertain or low-confidence outputs, it synthesizes a new, more accurate answer by jointly reasoning over their partial evidence. A detailed description of the Fusion agent’s implementation and prompt design is provided in Appendix A.4.
@@ 4.1 Experimental Setup
Evaluation Datasets and Metrics. We evaluate TableDART on seven diverse benchmarks across two tasks: Table Question Answering (TQA) and Table Fact Verification (TFV). For TQA, which requires interpreting complex queries (Ren et al., 2020; 2021) to perform grounded reasoning over tabular evidence, we use five benchmarks: WTQ (Pasupat and Liang, 2015), TABMWP (Lu et al., 2023), TAT-QA (Zhu et al., 2021), HiTab (Cheng et al., 2022), and FeTaQA (Nan et al., 2022). For TFV, we use TabFact (Chen et al., 2020) and InfoTabs (Gupta et al., 2020). Following the established protocol (Zheng et al., 2024; Liu et al., 2025), we use Accuracy for WTQ, TABMWP, TAT-QA, HiTab, TabFact, InfoTabs, while using BLEU score for FeTaQA as it is a generative free-form response task. Detailed statistics and further experimental setup details are provided in Appendix A.5.
Baselines. We compare TableDART against a comprehensive set of baselines, including (1) Table-as-Text models: Llama-2-7B (Touvron et al., 2023), Llama3-Instruct-8B (Dubey et al., 2024), TableLlama-7B (Zhang et al., 2024), and TableGPT2-7B (Su et al., 2024); (2) Table-as-Image models: MiniGPT-4-7B (Zhu et al., 2024a), mPLUG-Owl-7B (Ye et al., 2023), mPLUG-Owl2-7B (Ye et al., 2024), LLaVA v1.5-7B (Liu et al., 2024), Table-LLaVA-7B (Zheng et al., 2024), Qwen-VL-7B (Bai et al., 2023), InternLM-XComposer2-7B (Zhang et al., 2023), Monkey-7B (Li et al., 2024), TabPedia-7B (Zhao et al., 2024), SynTab-LLaVA-7B (Zhou et al., 2025), MiniCPM-V-2.6-8B (Yao et al., 2024), Qwen2.5-VL-7B (Bai et al., 2025), and Ovis2-8B (Lu et al., 2024); (3) Table-as-Multimodality models: HIPPO-8B (Liu et al., 2025) and Google Gemini 2.0 Flash (Comanici et al., 2025).
We employ TableGPT2-7B and Ovis2-8B as TableDART’s primary single-modality models. To demonstrate generalization, we additionally evaluate a variant by replacing the Table-as-Image model Ovis2-8B with Qwen2.5-VL-7B while retaining TableGPT2-7B. We use Google Gemini 2.0 Flash to implement the Fusion agent. Comparing TableDART against these base models directly demonstrates the effectiveness of our framework. Appendix A.5.2 provides a detailed rationale for model selection and clarifies the sourcing of baseline results adopted from previous studies. Notably, TableDART is a general framework that can be seamlessly integrated with a wide range of existing LLMs, VLMs and MLLMs.
@@ 4.2 Effectiveness Analysis
TableDART (TableGPT2-7B + Qwen2.5-VL-7B) 69.29 72.61 59.07 71.13 29.87 77.94 71.46 70.25
TableDART (TableGPT2-7B + Ovis2-8B) 70.58 84.54 62.05 74.37 36.11 81.37 76.22 74.86
State-of-the-Art Performance. Table 4.2 shows the comparison results of TableDART with baselines. From the results, we can observe that: (1) Compared to all single-modality based methods, TableDART consistently achieves superior or highly competitive results, including surpassing its constituent models like TableGPT2-7B and Ovis2-8B, demonstrating that the framework successfully combines their strengths to achieve a result neither could reach alone. (2) Compared to multi-modality based methods, TableDART’s dynamic routing mechanism proves more effective than the MLLM-based HIPPO, outperforming it on five of the seven benchmarks. (3) TableDART outperforms the powerful Fusion agent model (Google Gemini 2.0 Flash), confirming that the performance gains are driven by the intelligent routing mechanism, not just the capacity of the Fusion agent’s backbone. (4) TableDART demonstrates strong generalization across backbone models. When instantiated by replacing the Table-as-Image model with Qwen2.5-VL-7B while retaining TableGPT2-7B, the framework achieves a substantial +7.14% average accuracy gain over the single-modality Qwen2.5-VL baseline, validating its model-agnostic effectiveness. Furthermore, the ‘Average’ column summarizes the average performance, showing that TableDART achieves the strongest results among all baselines, surpassing the best multi-modality tabular model, HIPPO-8B, by a decisive +4.02% in average accuracy and +2.93 points in BLEU score.